OpenStudy (anonymous):

The graph of a limacon curve is given. Without using your graphing calculator, determine which equation is correct for the graph. a circle symmetric to the x axis where the majority of the circle lies in quadrants two and three [-5, 5] by [-5, 5] r = 2 - 2 cos θ r = 3 - 2 cos θ r = 3 - cos θ r = 3 - sin θ

OpenStudy (loser66):

To be honest, I cheat by using graphing tool to get the answer, it's \[r =3-2cos\theta\] algebraic way , \[r =2-2cos\theta\]has value r = 0 when theta = 0. and \[r=1-\sqrt2~when~\theta=\frac{\pi}{4}\] . the graph is not a circle.

OpenStudy (anonymous):

it's okay thanks anyway :)

OpenStudy (anonymous):

That answer is wrong.

OpenStudy (anonymous):

When I graph it I get r = 3 - cos θ?????

OpenStudy (anonymous):

r = 3 - cos θ 100%